# Decoding Noise in Nanofluidic Systems: Adsorption versus Diffusion Signatures in Power Spectra

### v1

@@ Abstract
Adsorption processes play a fundamental role in molecular transport through nanofluidic systems, but their signatures in measured signals are often hard to distinguish from other processes like diffusion. In this paper, we derive an expression for the power spectral density (PSD) of particle number fluctuations in a channel, accounting for diffusion and adsorption/desorption to a wall. Our model, validated by Brownian dynamics simulations, is set in a minimal yet adaptable geometry, allowing us to eliminate the effects of specific geometries. We identify distinct signatures in the PSD as a function of frequency $f$, including $1/f^{3/2}$ and $1/f^{1/2}$ scalings related to diffusive entrance and re-entrance effects, and a $1/f^{2}$ scaling associated with adsorption. These scalings appear in key predicted quantities – the total number of particles in the channel and the number of adsorbed or unadsorbed particles – and can dominate or combine in non-trivial ways depending on parameter values. Notably, when there is a separation of timescales between diffusion within the channel and adsorption/desorption times, the PSD can exhibit two distinct slopes in some of the predicted quantities. We provide a phase diagram to classify experimental systems based on predicted PSD shapes. These PSDs reflect measured properties in physical systems on the nano- and micro-scale, such as ion channels, nanopores, and electrochemical sensors, potentially offering insights into noisy experimental data.
@@ I Introduction
In particular, it is difficult to rationalize the coupled effects of adsorption and diffusive transport. Previous theoretical work has investigated the effect on the PSD of a single binding site in a cylindrical channel Berezhkovskii, Pustovoit, and Bezrukov (2002) and adsorption to the walls of a pore in combination with the effect of reservoirs Gravelle, Netz, and Bocquet (2019); Robin et al. (2023a). In these models, it can be difficult to distinguish the effects of adsorption from other model-specific features affecting particle dynamics, especially the geometry, which introduces several time- and length-scales Marbach (2021); Robin et al. (2023a). This has led some authors to linearly sum PSDs induced by uniquely diffusion or adsorption to account for experimental data Zevenbergen et al. (2009), a likely flawed attempt in systems where the timescales for both phenomena can be similar Knowles et al. (2021); Gadgil et al. (2004). In particular, in experimental data involving adsorption and diffusion, it is still unclear when adsorption yields specific signatures in the PSD, as it is sometimes apparently linked to the emergence of two distinct slopes Zevenbergen et al. (2009); Robin et al. (2023a) and sometimes not Knowles et al. (2021); Secchi et al. (2016).
To answer this question, we introduce a simple framework to explore and separate the effects of diffusion and adsorption on power spectra. Our model involves non-interacting particles diffusing along a one-dimensional line, with a segment where particles can adsorb and become immobile – see Fig. 2. This model is versatile and can easily be adapted to various geometries, such as channels Marbach (2021) or electrochemical devices Zevenbergen et al. (2009). Its simplicity allows us to isolate geometric effects from those of diffusion and adsorption. We derive analytical expressions, validated by simulations, for the auto-correlation function and PSD of particle number fluctuations for all particles in the channel, as well as for free (unadsorbed) and bound (adsorbed) particles (Sec. II). The resulting PSDs are complex combinations of the purely diffusive and adsorbing processes, showing distinct scalings like $1/f^{1/2}$, $1/f^{3/2}$, and $1/f^{2}$, which combine non-linearly. We rationalize the mechanisms behind each of these scaling laws: diffusive re-entrance and entrance effects and adsorption/desorption phenomena, respectively. Crucially, when there is significant difference in timescales between diffusion across the channel and adsorption/desorption, the PSDs of some but not all of the predicted quantities exhibit two distinct slopes (Sec. III). Finally, we discuss strategies to solve the inverse problem of determining adsorption properties from experimental data (Sec. IV).
@@ II.1 Setup to Probe the Effect of Adsorption
In choosing such a general 1D system rather than a specific geometry, our model is easily adaptable to different geometries. For example, it can be applied to systems with two binding regions on either side of the domain, as in many electrochemical cells, or a single binding region on one side, as in test assays. We can also change $L$ and $L_{0}$ at will to resemble systems where the binding region is large ($L\sim L_{0}$, as in electrochemical devices and assays) or small ($L\gg L_{0}$, as in many nanopore systems). Some of these different geometries are illustrated in Fig. 1.
@@
where $s$ is the Laplace frequency, $m_{0}$ and $m_{1}$ are functions of $s$, and $A_{C}=B_{C}$ is a constant determined from the boundary conditions. Note that the expression for $\hat{p}$ is only true inside the channel where $|x|\leq L_{0}/2$.
@@ II.2 Number Fluctuations in the Channel
[…]
@@ II.3 Spectral Density
[…]
@@
Note that $s=2i\pi f$ in the above expression. Eq. (12) is the main theoretical result of this paper. In its current form, it is hardly amenable to physical interpretation, however, and our goal in the following sections is to uncover limiting regimes and mechanistic behaviour from this expression.
@@ III Results
In spite of its apparent simplicity, our system is governed by four key timescales:
@@
These timescales can be estimated for many nanotechnological systems and we report orders of magnitude in Table 1. Remarkably, there is no standard case, and relative magnitudes can vary strongly from one system to another.
@@
In the following, we start by investigating limit cases of pure diffusion or only adsorption. This will allow us to establish a few canonical fluctuation regimes before turning to the full, much more complex system.
Related to this, we work in the following reduced units. In all cases the characteristic length scale is $\ell_{0}=L_{0}$, and the characteristic timescale is $\tau_{0}=\tau_{\rm cha}=L_{0}^{2}/2D$, except in the case of pure adsorption/desorption where $D=0$, in which case we use $\tau_{0}=1/f_{0}$, where $f_{0}=(k_{\rm on}+k_{\rm off})/2\pi$ and represents the corner frequency in the PSD. We compare the analytical results, derived above, to particle-based simulations of the system (see Appendix C for details on computational methods).
@@ III.1 Pure Diffusion
The theoretical (line) and simulation (dots) spectra for the total number of particles are shown in Fig. 3 (purple) and are in perfect agreement. The spectrum clearly involves multiple regimes; more specifically, a plateau at low frequencies and then two scaling regimes with exponents $-1/2$ and $-3/2$. The scaling regimes can be interrogated by considering the limits of Eq. (13).
[…]
@@
with a scaling as $1/f^{3/2}$. Between these two regimes, the slope of the PSD appears to vary steadily between the two limits; however, this masks a well-defined underlying scaling. To find this, we take $L\rightarrow\infty$ and $f\rightarrow 0$
@@
predicting a scaling as $1/f^{1/2}$. Physically, we can interpret these different regimes as follows.
@@
Remembering that $p_{0}=1/L$ and taking $L\gg L_{0}$ for simplicity, we find $\tau_{\rm corr}=\frac{L_{0}^{2}}{3D}=2\tau_{\rm cha}/3$. The plateau amplitude thus depends on the mean time particles spend inside the channel. It is interesting that the amplitude does not depend on system size $L$ but actually probes the channel size $L_{0}$, and the only dependence of the plateau on $L$ is the frequency at which it begins.
High frequency $1/f^{3/2}$ is linked to diffusive particle crossings in and out of the channel.
At high frequencies, corresponding to timescales shorter than the time for a particle to diffuse across the channel, $\tau_{\rm cha}$, we see a characteristic $1/f^{3/2}$ scaling. This has been observed in previous works Marbach (2021); Lax and Mengert (1960); Zevenbergen et al. (2009); Berezhkovskii, Pustovoit, and Bezrukov (2002); Minh, Rotenberg, and Marbach (2023). Its origin is particles crossing from one region to another (i.e. from channel to reservoir and vice versa), by diffusion. This result is independent of the system geometry.
Intermediate $1/f^{1/2}$ is linked to re-entrance effects.
The intermediate regime between these two timescales – where we observe a slope in the PSD lower than $3/2$, but have not yet reached a plateau – is a signature of the underlying $1/f^{1/2}$ scaling. While its presence has been highlighted in Ref. [Marbach, 2021], its origin has not been identified. Here, we show this scaling arises from re-entrance effects, i.e. particles diffusing back into the channel that have previously diffused out of the channel at intermediate timescales. This leads to long-lived correlations that produce an additional scaling regime in the PSD before correlations are removed over timescales longer than $\tau_{\rm sys}$.
As further evidence of this, we use simulations to probe a modified version of our model system as defined in Fig. 2. Here, the particles diffuse normally along the line $P$, up to the point that they cross the channel boundaries. When a particle leaves the channel, however, it is re-allocated to a location randomly picked in the reservoir. In this way, we artificially remove the “return” correlation in particle trajectories. Indeed, this entirely removes the scaling regime as $1/f^{1/2}$ between the channel and system timescales, as shown in Fig. 3 (green dots). As a result, the onset of the plateau regime occurs around a frequency corresponding to the time to diffuse across the channel $1/\tau_{\rm cha}$, and not at $1/\tau_{\rm sys}$; by removing the correlations when particles exit the channel, they are no longer probing the finite system size, but only the finite channel size.
@@ III.2 Only Adsorption/Desorption
where $\mathcal{N}_{\rm off}=\mathcal{N}L_{0}p_{0}$ is the mean number of free particles, and we recall that $f_{0}=2\pi/(k_{\rm off}+k_{\rm on})$. This expression is consistent with works which investigate exclusively adsorption/desorption processes Zevenbergen et al. (2009); Nestorovich et al. (2002).
@@
Figure 4: Power spectral density of number fluctuations with no diffusion and only adsorption/desorption processes. Spectrum for fluctuations in the number of unadsorbed particles is shown, which is equivalent to that for adsorbed particles. Dashed lines indicate $\tau_{\rm ads}=1/k_{\rm on}$ and $\tau_{\rm des}=1/k_{\rm off}$. A solid line indicates $f_{0}$, the corner frequency, which is the mean of the two rates $f_{0}=(k_{\rm on}+k_{\rm off})/2$. There is a clear $1/f^{2}$ scaling at high frequencies and a plateau below $f_{0}$. In this case, $L_{0}=10\ell_{0}$, $k_{\rm on}=0.009f_{0}$, and $k_{\rm off}=2f_{0}$. We use separated adsorption and desorption rates to demonstrate the dependence of the corner primarily on the faster rate (here, $k_{\rm off}$).
The shape of the spectrum is a Lorentzian, scaling as $1/f^{2}$ at high frequencies. This can be intuitively understood as, at very short timescales, the fluctuating signal appears as individual step functions. Transforming into frequency space, a step function becomes $1/f$, which is then squared to obtain the PSD and $1/f^{2}$ scaling. At low frequencies, or long timescales, the fluctuating signal is “smeared out” and resembles a white noise, or a plateau in the PSD. Remarkably, the binding/unbinding process corresponds to a single shot noise process, which again intuitively makes sense because binding and unbinding events are uncorrelated and thus follow a Poisson distribution.
@@
Similar noise processes and therefore spectral signatures are observed in various other physical systems. For example, this noise is similar to random telegraph or “popcorn” noise – noise observed in semiconductors where there are sudden transitions between two (or sometimes more) current levels Machlup (1954). In semiconductors, this is often due to the generation and trapping of charge carriers. Analogously, here, this noise is due to particles having only two states, “on” or adsorbed, and “off” or unadsorbed. Equally, the process can be mapped to scenarios where noise arises from signals involving “free particle pulses”. Here, the “pulse” is the mean number of free particles $\mathcal{N}_{\rm off}$ and occurs with mean frequency $k_{\rm on}$, which corresponds to the mean time particles stay free. Considering that $\mathcal{N}_{\rm off}k_{\rm on}=\mathcal{N}_{\rm on}k_{\rm off}$, the process can also be interpreted as “bound particle pulses”. The subtlety is that the amplitude of each of these pulses decays exponentially with a timescale which is the mixture of the two timescales at play $\tau_{0}$. The exponential distribution of the pulses is a specificity of this system, and thus, compared to other processes whose length in time are constant, as in Ref. [Knowles, Mackay, and Thorneywork, 2024], the shape of the PSD has no oscillations and is a true Lorentzian Verveen and DeFelice (1974).
Figure 5: Slow diffusion, fast adsorption/desorption: Case 1 – $\tau_{\rm cha}\gg\tau_{\rm ads}\sim\tau_{\rm des}$. Left: total PSD. Middle: bound PSD. Right: free PSD. We use $L=2\ell_{0}$, $k_{\rm on}=20/\tau_{0}$, and $k_{\rm off}=60/\tau_{0}$. Note here that $k_{\rm on}\leq k_{\rm off}$. The dotted curve indicates the pure diffusive case with the same parameters (but $k_{\rm on}=0$). Similarly, the dash-dotted curve indicates the pure adsorption/desorption case with the same parameters (but $D=0$). Relevant timescales are labeled and marked with dashed vertical lines.

### v2

@@ Abstract
Adsorption processes play a fundamental role in molecular transport through nanofluidic systems, but their signatures in measured signals are often hard to distinguish from other processes like diffusion. In this paper, we derive an expression for the power spectral density (PSD) of particle number fluctuations in a channel, accounting for diffusion and adsorption/desorption to a wall. Our model, validated by Brownian dynamics simulations, is set in a minimal but adaptable geometry, allowing us to eliminate the effects of specific geometries. We identify distinct signatures in the PSD as a function of frequency $f$, including a $1/f^{3/2}$ scaling related to diffusive entrance/exit effects, and a $1/f^{2}$ scaling associated with adsorption. These scalings appear in key predicted quantities – the total number of particles in the channel and the number of adsorbed or unadsorbed particles – and can dominate or combine in non-trivial ways depending on parameter values. Notably, when there is a separation of timescales between diffusion inside the channel and adsorption/desorption times, the PSD can exhibit two distinct corners with well-separated slopes in some of the predicted quantities. We provide a strategy to identify adsorption and diffusion mechanisms in the shape of the PSD of experimental systems on the nano- and micro-scale, such as ion channels, nanopores, and electrochemical sensors, potentially offering insights into noisy experimental data.
@@ I Introduction
In particular, it is difficult to rationalize the coupled effects of adsorption and diffusive transport. Previous theoretical work has investigated the effect on the PSD of a single binding site in a cylindrical channel Berezhkovskii, Pustovoit, and Bezrukov (2002) and adsorption to the walls of a pore in combination with the effect of reservoirs Gravelle, Netz, and Bocquet (2019); Robin et al. (2023a). In these models, it can be difficult to distinguish the effects of adsorption from other model-specific features affecting particle dynamics, especially the geometry, which introduces several time- and length-scales Marbach (2021); Robin et al. (2023a). This has led some authors to linearly sum PSDs induced by uniquely diffusion or adsorption to account for experimental data Zevenbergen et al. (2009), a likely flawed attempt in systems where the timescales for both phenomena can be similar Knowles et al. (2021); Gadgil et al. (2004). In particular, in experimental data involving adsorption and diffusion, it is still unclear when adsorption yields specific signatures in the PSD, as it is sometimes apparently linked to the emergence of two distinct slopes Zevenbergen et al. (2009); Robin et al. (2023a) and sometimes only one slope Knowles et al. (2021); Secchi et al. (2016).
To answer this question, we introduce a simple framework to explore and separate the effects of diffusion and adsorption on power spectra. Our model involves non-interacting particles diffusing along a one-dimensional line, with a segment where particles can adsorb and become immobile – see Fig. 2. Although the model is only an approximation to real physical systems, its simplicity is useful in two ways. Firstly, it allows us to isolate geometric effects from those of diffusion and adsorption within the channel, since in 1D the channel and reservoirs have only two length scales and therefore only two diffusive timescales. Secondly, it allows us to derive exact analytical expressions, validated by simulations, for the auto-correlation function and PSD of particle number fluctuations for all particles in the channel, as well as for free (unadsorbed) and bound (adsorbed) particles (Sec. II). The resulting PSDs are complex combinations of the purely diffusive and adsorbing processes, showing distinct scalings, like $f^{-3/2}$ and $f^{-2}$, which combine non-linearly. We rationalize the mechanisms behind each of these scaling laws: diffusive entrance/exit effects and adsorption/desorption phenomena, respectively. Crucially, when there is significant difference in timescales between diffusion across the channel and adsorption/desorption, the PSDs of some – but not all – of the predicted quantities exhibit two distinct slopes (Sec. III). Finally, we discuss strategies to solve the inverse problem of determining adsorption properties from experimental data (Sec. IV).
@@ II.1 Setup to Probe the Effect of Adsorption
There are other possible approaches to include the effect of reservoirs in our system; for example, another line where particles can “hop out” of the system for $\lvert x\rvert>L_{0}/2$ (as in Ref. [Marbach, 2021]). However, as the impact of the dimensions and geometry of reservoirs has been studied previously Robin et al. (2023a); Gravelle, Netz, and Bocquet (2019) we focus on this 1D representation.
In choosing such a general 1D system rather than a specific geometry, we reduce the number of length scales in the system to two, $L$ and $L_{0}$. This reduces the number of corresponding diffusive timescales in the system to two, the minimum number possible in a system containing a channel and reservoirs. While more complicated 3D geometries may be more accurate representations of real systems, the additional length and timescales introduced would add complexities to the power spectra. Furthermore, even if the system is 3D, effective 1D behavior can still occur where particles are diffusing in a region with one length scale much shorter than the other(s) – e.g. in nanopores Marbach (2021). Even with only two length scales, we can change $L$ and $L_{0}$ at will to resemble systems where the adsorption region is large ($L\sim L_{0}$, as in electrochemical devices and assays) or small ($L\gg L_{0}$, as in many nanopore systems). Lastly, our 1D model allows us to solve for the predicted quantities exactly and gain physical insights.
@@
where $s$ is the Laplace frequency, $m_{0}(s)=\sqrt{s/D}$ and $m_{1}(s)=\sqrt{s(s+k_{\rm off}+k_{\rm on})/D(s+k_{\rm off})}$, and $A_{C}=B_{C}$ is a constant determined from the boundary conditions. Note that the expression for $\hat{p}$ is only true inside the channel where $|x|\leq L_{0}/2$.
@@ II.2 Number Fluctuations in the Channel
$\partial_{t}C_{N}(t)$ $=\mathcal{N}\int_{-L_{0}/2}^{L_{0}/2}\left(\partial_{t}p(x,t)+\partial_{t}q(x,t)\right)dx$ (7)
@@ II.3 Spectral Density
$S_{N}(f)=\int_{-\infty}^{\infty}C_{N}(t)e^{-2\pi ift}dt.$ (9)
The correlation function for a stationary process like ours is even ($S_{N}(f)=2\int_{0}^{\infty}C_{N}(t)e^{-2\pi ift}dt$), so using the Laplace transform
@@
Note that $s=2i\pi f$ in the above expression. Eq. (12) is the main theoretical result of this paper. In its current form, it is hardly amenable to physical interpretation, however, and our goal in the following sections is to uncover limiting regimes and mechanistic behavior from this expression.
@@ III Results
Our system is governed by four key timescales:
@@
These timescales can be estimated for many nano-technological systems and we report orders of magnitude in Table 1. Remarkably, there is no standard case, and relative magnitudes can vary strongly from one system to another. Our goal is not to imitate these systems exactly with our model, but rather to illustrate the diversity of behavior which can be observed in power spectra with only two physical processes.
@@
In the following, we start by investigating limit cases of pure diffusion or only adsorption. While the scaling laws associated with these processes in isolation are known, the full problem of interest here involves a non-trivial combination of these two canonical mechanisms. As such, we provide a detailed explanation of how scalings emerge from fluctuations associated with these separate phenomena as context for later results. Scaling laws for adsorption and diffusion are also accessible experimentally and provide a useful reference point for experimentalists.
Related to this, we work in the following reduced units. In all cases the characteristic length scale is $\ell_{0}=L_{0}$, and the characteristic timescale is $\tau_{0}=\tau_{\rm cha}=L_{0}^{2}/2D$, except in the case of pure adsorption/desorption where $D=0$, in which case we use $\tau_{0}=1/f_{0}$, where $f_{0}=(k_{\rm on}+k_{\rm off})/2\pi$ and represents the corner frequency in the PSD. We compare the analytical results, derived above, to particle-based simulations of the system (see Appendix C for details on computational methods). Again, we generally choose a 1D geometry for our simulations to provide a robust verification of our analytical results.
@@ III.1 Pure Diffusion
The theoretical (line) and simulation (dots) spectra for the total number of particles are shown in Fig. 3 (purple) and are in perfect agreement. The spectrum clearly involves multiple regimes; more specifically, a plateau at low frequencies and then two scaling regimes with exponents 1/2-1/2 and 3/2-3/2. The scaling regimes can be interrogated by considering the limits of Eq. (13).
Figure 3: Power spectral density of number fluctuations for the purely diffusive case, with $L=20\ell_{0}$. Dashed lines indicate $\tau_{\rm sys}=L^{2}/2D$ and $\tau_{\rm cha}=L_{0}^{2}/2D$. Expected scalings for high and intermediate frequency regimes of $f^{-3/2}$ and $f^{-1/2}$ respectively are indicated. Data in purple shows the spectrum for the total number of particles from simulations. Data in green shows the corresponding spectrum for simulations in 3D, with the same parameters. The 3D result plateaus for frequencies smaller than $1/\tau_{\rm cha}=2D/L_{0}^{2}$.
@@
with a scaling as $f^{-3/2}$. Between these two regimes, the slope of the PSD appears to vary steadily between the two limits; however, this masks a well-defined underlying scaling. To find this, we take $L\rightarrow\infty$ and $f\rightarrow 0$
@@
predicting a scaling as $f^{-1/2}$. Physically, we can interpret these different regimes as follows.
@@
Remembering that $p_{0}=1/L$ and taking $L\gg L_{0}$ for simplicity, we find $\tau_{\rm corr}=L_{0}^{2}/3D=2\tau_{\rm cha}/3$. The plateau amplitude thus depends on the mean time particles spend inside the channel. It is interesting that the amplitude does not depend on system size $L$ but actually probes the channel size $L_{0}$, and the only dependence of the plateau on $L$ is the frequency at which it begins.
High frequency $f^{-3/2}$ is linked to diffusive particle crossings in and out of the channel.
At high frequencies, corresponding to timescales shorter than the time for a particle to diffuse across the channel, $\tau_{\rm cha}$, we see a characteristic $f^{-3/2}$ scaling Marbach (2021); Lax and Mengert (1960); Zevenbergen et al. (2009); Berezhkovskii, Pustovoit, and Bezrukov (2002); Minh, Rotenberg, and Marbach (2023). Its origin is particles crossing from one region to another (i.e. from channel to reservoir and vice versa), by diffusion. These events correspond to short time scales and are essentially independent of the system geometry.
Intermediate $f^{-1/2}$ is linked to re-entrance effects.
The intermediate regime between these two timescales – where we observe a slope in the PSD lower than $3/2$, but have not yet reached a plateau – is a signature of a $f^{-1/2}$ scaling Marbach (2021); MacFarlane (1950); Burgess (1953); Van Vliet and Van der Ziel (1958); Lax and Mengert (1960); Voss and Clarke (1976). This scaling arises from re-entrance effects, i.e. it emerges directly from the fact that particles in 1D always return to the origin. This intermediate scaling depends on dimensionality. In an $n$-dimensional system, the probability of particle returns goes as $1/t^{n/2}$, so in 1D the spectral scaling is $f^{-1/2}$, in 2D it is $\rm{ln}(1/\textit{f})$, and in 3D it is $1/f^{0}$, i.e. a plateau Voss and Clarke (1976); MacFarlane (1950); Burgess (1953). The cross-over from 1D to 3D is the point at which particles no longer inevitably return to their origin, so the re-entrance correlations disappear, as shown in Fig. 3 (green dots). In more realistic designs, and even in 3D systems, there may be effective 1D or 2D scaling behavior if one or more dimension is small relative to the others Voss and Clarke (1976). In our 1D model, a plateau below $1/\tau_{\rm sys}$ appears due to finite system size. These timescales may be inaccessible in experimental systems. Overall, these low frequency scalings are linked to the reservoir geometry, which is not our focus here.
@@ III.2 Only Adsorption/Desorption
where $\mathcal{N}_{\rm off}=\mathcal{N}L_{0}p_{0}$ is the mean number of free particles, and we recall that $f_{0}=(k_{\rm off}+k_{\rm on})/2\pi$. This is consistent with works which investigate exclusively adsorption/desorption processes Zevenbergen et al. (2009); Nestorovich et al. (2002).
@@
Figure 4: Power spectral density of number fluctuations with no diffusion and only adsorption/desorption processes. Spectrum for fluctuations in the number of unadsorbed particles is shown, which is equivalent to that for adsorbed particles. Dashed lines indicate $\tau_{\rm ads}=1/k_{\rm on}$ and $\tau_{\rm des}=1/k_{\rm off}$. A solid line indicates $f_{0}$, the corner frequency, which is the mean of the two rates $f_{0}=(k_{\rm on}+k_{\rm off})/2\pi$. There is a clear $1/f^{2}$ scaling at high frequencies and a plateau below $f_{0}$. In this case, $L=L_{0}=\ell_{0}$, $k_{\rm on}=0.027f_{0}$, and $k_{\rm off}=6.3f_{0}$. We use separated adsorption and desorption rates to demonstrate the dependence of the corner primarily on the faster rate (here, $k_{\rm off}$).
The shape of the spectrum is a Lorentzian, scaling as $1/f^{2}$ at high frequencies. This can be intuitively understood as, at very short timescales, the fluctuating signal appears as individual step functions. Transforming into frequency space, a step function becomes $1/f$, which is then squared to obtain the PSD and $1/f^{2}$ scaling. At low frequencies, or long timescales, the fluctuating signal is “smeared out” and resembles a white noise, or a plateau in the PSD. The binding/unbinding process corresponds to a single shot noise process, which again intuitively makes sense because binding and unbinding events are uncorrelated and thus follow a Poisson distribution.
@@
Figure 5: Slow diffusion, fast adsorption/desorption: Case 1 – $\tau_{\rm cha}\gg\tau_{\rm ads}\sim\tau_{\rm des}$. Left: total PSD. Middle: bound PSD. Right: free PSD. We use $L=2\ell_{0}$, $k_{\rm on}=125/\tau_{0}$, and $k_{\rm off}=375/\tau_{0}$. Note here that $k_{\rm on}\leq k_{\rm off}$. The dotted curve indicates the pure diffusive case with the same parameters (but $k_{\rm on}=0$). Similarly, the dash-dotted curve indicates the pure adsorption/desorption case with the same parameters (but $D=0$). Relevant timescales are labeled and marked with dashed vertical lines.
Similar noise processes and therefore spectral signatures are observed in various other physical systems. For example, this noise is similar to random telegraph or “popcorn” noise – noise observed in semiconductors where there are sudden transitions between two (or sometimes more) current levels Machlup (1954). Analogously, here, this noise is due to particles having only two states, “on” or adsorbed, and “off” or unadsorbed. Equally, the process can be mapped to scenarios where noise arises from signals involving “free particle pulses”. Here, the “pulse” is the mean number of free particles $\mathcal{N}_{\rm off}$ and occurs with mean frequency $k_{\rm on}$, which corresponds to the mean time particles stay free. Considering that $\mathcal{N}_{\rm off}k_{\rm on}=\mathcal{N}_{\rm on}k_{\rm off}$, the process can also be interpreted as “bound particle pulses”. The subtlety is that the amplitude of each of these pulses decays exponentially with a timescale which is the mixture of the two timescales at play $\tau_{0}$. The exponential distribution of the pulses is a specificity of this system, and thus, compared to other processes whose length in time are constant, as in Ref. [Knowles, Mackay, and Thorneywork, 2024], the shape of the PSD has no oscillations and is a true Lorentzian Verveen and DeFelice (1974).